Capturing the Particulars of Classroom Practice Kevin Karplus

31 May 1994, revised 24 December 1994

Board of Studies in Computer Engineering University of California, Santa Cruz Santa Cruz, CA 95064

This essay was prompted by an exercise suggested by the American Association for Higher Education, which asked for a “telling episode or some incident of classroom practice that reveals something distinctive” and provided three possible strategies to document the episode or incident: videotape, classroom visit, or case report.

Unfortunately, “telling episodes” do not happen on demand and to schedule, so neither of the first two strategies seems appropriate, though either is less work than writing a case report. If the goal of the exercise had been to observe a typical class, rather than a “telling episode”, I would probably have chosen to use videotape.

1 Course: Art of the Book in the Computer Age

The particular incident I wish to write about occurred in Spring 1994, in the class The Art of the Book in the Computer Age. The course is a “topics” course, which means that students can get credit towards their general education requirements by taking the course. There are no prerequisite courses, though I do ask that students have some familiarity with computers—at least to the point where they can learn to use a new Macintosh program and edit files on a Unix system primarily on their own.

The course attempts to bridge the cultural gaps between the sciences, the arts, and the humanities while teaching a marketable skill in desktop publishing. As a result there is a wide range of backgrounds, motivation, and academic culture for the students in the class—very different from the fairly homogeneous mix in most engineering classes.

The classroom presentations include many styles that are unusual for an engineering professor, including 35mm slide lectures, workshops in which the students critique each others’ work, and classroom discussions analyzing work done by others, as well as more familiar lecture/discussions using the blackboard or overhead projector. Although I created this class, and this was the fourth time I taught it, I am still not entirely comfortable teaching what is primarily an art class with no background in the visual arts.

2 Episode: Discussion on Poster Design

In May, one of the art galleries in Porter College decided to discard their collection of posters announcing various shows at the gallery. This huge pile of different posters all designed with a similar goal gave me the idea for another lecture or discussion in the course, comparing the different approaches for attracting people to an art show. I picked up about twenty-five different posters on Wednesday (25 May) to use in Friday’s class.

I usually do not cover much poster design in the course (other than a slide lecture or two on poetry broadsides from the library’s Special Collections, which I had already done that quarter), but a discussion on poster design was well within the range of things that are suitable for the class. Because these posters were a spur-of-the-moment addition to the course, I did not have time to do a careful analysis of them, nor to construct a coherent framework for discussing them.

The most common format for the lecture/discussions in the course this quarter has been for me to display one piece at a time to the class using slides, overhead transparencies, or, occasionally, the piece itself, then to use a Socratic dialog to elicit useful comments from the class. This approach had been fairly successful spring quarter, with most of the students who attended class participating in the discussion. However, as is usually the case with class discussions, two or three of the students generally contributed much more than the rest.

As an experiment, I decided to try a different format for the discussion—one that relied much less on me as a guide and that would allow everyone to participate more or less equally in the discussion.

I brought the stack of posters to class and spread them out on the table in front of the class, instructing each student to take a poster and to make notes about the poster for 5–10 minutes. The students were told to pick a poster they liked, one they disliked, or just one they thought they could say something about. We then arranged the seats in a circle and went around the room with each student having up to five minutes to present his or her observations about the poster. If the student presenting the poster had less to say, then the discussion was opened up for everyone else to add comments.

3 Reflections

This format for the discussion was productive in several ways:

• Everyone got roughly equal time to speak. In fact, since some of the shyer students arrived on time, and some of the more active participants were late, the shy students may have had more to say about their selected posters than the normally more aggressive students.

• I chose the order of student presentations to be roughly the order in which students arrived in the classroom, so that those who had had longest to study their posters got to speak first. I made the order clear to the students, so they would know about when it would be their turn.

• Individuals had a sense of ownership in presenting “their” poster to the class, but were free to criticize it or praise it without ego involvement (unlike the workshop we’d had earlier in the quarter where students discussed each other’s work).

• Most of the students presented fairly detailed critiques of their posters, providing much the same sort of analysis that I would have hoped to get from a more directed discussion. To a large extent, I believe this is a result of the large number of directed discussions we’ve had during the quarter.

• One student did a comparative analysis of two posters that used similar design techniques (black background, white-on-black text, and a single graphic element from the art show), but with strikingly different effect. This comparative analysis was not something I had planned for the discussion, but it was a good addition. Another comparative analysis came up because two students had chosen posters that were clearly designed at the same time by the same designer, for two different shows that were on at the same time. It was interesting for the students to notice how small changes (like the background color of the graphic image or the font used for the title line) made an enormous difference in the effect of almost identically designed posters. Again, this comparative analysis was initiated by the students, not forced by any guidance on my part.

• One discussion, which I allowed to continue past the five-minute cutoff, was of a particularly unsuccessful poster, whose underlying organization (as a calendar) was not visible from a distance, but only close up. I did direct the students slightly on this discussion, asking them how they would make the calendar more visible as an organizing theme—several good ideas were suggested.

There were some negative aspects to the discussion format:

• Many of the students have gotten into the habit of trickling in late to class. Because students kept arriving, I extended the initial note-taking time from ten minutes to over twenty minutes. This was not too long (the early students were still busy at the end of the time), but it did cut into the discussion time.

• Because I let one of the discussions that was particularly productive run over the five-minute limit I had set, the class did not end on time. I gave students permission to leave when the normal class time was up, but everyone stayed to hear the last presentation. The final presentation was a bit hurried as a result of the overtime.

• The five-minute-per-student format only worked because so few students show up for class on Friday (the class meets 5 p.m.–6:10 p.m., and many students decide to start their weekend early). I would have difficulty using this format on a day when attendance is heavier—it might also be more difficult to get all the students to participate productively when the less-dedicated students are present.

• The other students had to be able to see the poster being discussed as the student pointed out the things he or she had observed. This was solved by having an adjacent student hold up the poster, but it would be difficult to use this approach to sharing the image for any objects smaller than these large posters.

• Posters read quite differently from a distance than they do close up, and most of the students had only looked at their posters close up, sitting at their desks. Several of the comments made by students other than the presenter had to do with this difference in viewing distance.

I used the poster discussion again Fall 1994 with a much larger class (about 35 were present). I had originally intended to do the discussion near the end of the quarter, shortly after the poetry-broadside slide lecture/discussions. Because the discussions at the beginning of the quarter on letterheads and business cards had been greatly expanded this quarter, the poetry broadsides were delayed to the last week, and the poster discussion occurred on the last day of class.

The discussion format was still productive, but a lot of the students got a bit silly, since they had just finished turning in their final projects. It would probably be better to work the discussion in earlier next year. I still did not come up with a good way for the students to view the posters from a variety of distances—though the difference in how the posters read from different distances certainly did come out in the discussion.

I’d like to devise some other discussions that can be run this way, with students each presenting one example. I haven’t figured out how I’ll handle the problem of sharing the image while it is being discussed, particularly for some of the smaller images we discuss (business cards, for example). The only thing I can think of is to pre-prepare a set of slides, and give each student one to work on ahead of time.